# Certified Randomness implies Secure Classical Position-Verification

### v2

@@ 3 CVPV from Single-Round Certified Randomness
Completeness.
Completeness follows by completeness of $\mathscr{P}$.
@@
Let $\mathscr{P}=(V,P)$ be a one-round PoQ scheme that satisfies Definition 3. Then, 1 is a CVPV scheme that is sound in the quantum random oracle model.
@@
  • The guesser $Q$ holds register ${\color[rgb]{1,.5,0}E}={\color[rgb]{1,.5,0}B}$. She eavesdrops $\mathsf{ch}$ and runs $\mathcal{B}(t\geq 0)$ to output $\mathsf{ans}^{\prime}$.
With probability $p_{4}$, $\mathsf{ans}=\mathsf{ans}^{\prime}$ and $\mathsf{ans}$ is accepted by the verifier (denoted by event $ACC$). Thus, we have
$p_{4}\leq\Pr\left[\mathsf{ans}=\mathsf{ans}^{\prime}\;|\;ACC\right]\leq 2^{-H_{\min}\left(\mathsf{ans}\;|\;\mathsf{ch},{\color[rgb]{1,.5,0}E}\right)}$
$\implies$ $H_{\min}\left(\mathsf{ans}\;|\;\mathsf{ch},{\color[rgb]{1,.5,0}E}\right)\leq\log_{2}(1/p_{4})\leq O(\log\lambda),$
which violates certified randomness (Definition 3). ∎
@@ 4 CVPV from Multi-Round Certified Randomness
If the underlying PoQ-based certified randomness protocol uses more than round, we can naturally generalize our compiler just by composing our single-round compiler sequentially. We first give a natural way to do this in Section 4.1. This requires a protocol satisfying the stronger Definition 5 of certified randomness with sequential decomposition, but we show that any protocol satisfying the definition of certified randomness with single round entropy must satisfy the sequential decomposition property in Appendix A, lending its applicability to a wide range of possible instantiations. We then give a more clever way in Section 4.3 and Section 4.4 which is superior in the idealized model at the cost of practical robustness.
@@
An $\ell$-round PoQ protocol $\mathscr{P}$ is said to have sequential decomposition property if no pair of a QPT prover $P$ and an unbounded guesser $Q$ can succeed in the following security game with non-negligible probability:
@@
We also consider a weaker variant of Definition 5 by restricting the communication round between $P$ and $Q$, with the difference being highlighted:
@@
An $\ell$-round PoQ protocol $\mathscr{P}$ is said to have sequential certified randomness with restricted communication if no pair of a QPT prover $P$ and an unbounded guesser $Q$ can succeed in the following security game with non-negligible probability:
@@
For single-round $\mathscr{P}$, Definitions 3, 7, 5 and 6 all coincide.
We say that two spatially separated parties $(\mathcal{A},\mathcal{B})$ perform one round of simultaneous communication at time $t$ if $\mathcal{A}$ sends one (classical or quantum) message to $\mathcal{B}$ at time $t$ and vice versa.
@@ 4.1 Sequential Compiler
Let $\mathscr{P}=(V,P)$ be a PoQ scheme that satisfies Definition 6. Then, 2 is a secure CVPV scheme.
@@
Follows directly from Definition 6. $\mathcal{A}$ plays the role of the prover and $\mathcal{B}$ that of the eavesdropper. ∎
@@ 4.2 Sequential Decomposition from Repetition
This section uses the same notation as Appendix A. Although other multi-round approaches are possible, the most general sequential compilation approach is desirable due to practical robustness considerations. Section 4.1 proves that one can construct CVPV protocols from certified randomness protocols with the sequential decomposition property. However, the usual notion of certified randomness (Definition 3) is defined in terms of entropy lower bounds, but we cannot lower bound the CVPV protocol entropy conditioned on the guesser’s side information since she can communication with the prover. Therefore, it is not immediately obvious whether a certified randomness protocol defined with entropy is a sequential certified randomness protocol (Definition 5) defined by the guessing probability. We show that for a large class of multi-round certified randomness protocols, namely those that rely on repetition of single rounds with entropy bounds ‘on average’, all have the sequential decomposition property.
@@ A PoQ protocol $\mathscr{P}$ is said to be certified randomness from repetition if
To show that protocols satisfying Definition 7 can be used to construct CVPV protocols (illustrated in Fig. 8(b)) with the sequential decomposition property, we first consider an optimal adversary with the largest overall acceptance probability $\Pr[\Omega]$, where $\Omega$ denotes the event where the answers pass the protocol statistical test and all the guesses are correct. We denote the quantum channel of the $n$-round optimal adversary as $\mathcal{M}_{n}^{{}^{\prime}*}\circ\cdots\circ\mathcal{M}_{1}^{{}^{\prime}*}$. We then construct a modified adversary ¯$\bar{\mathcal{M}}_{n}^{{}^{\prime}}\circ\cdots\circ\bar{\mathcal{M}}_{1}^{{}^{\prime}}$ from this optimal adversary, and show that the modified adversary has the same $\Pr[\Omega]$. Then, we show the modified adversary also satisfies the non-signalling condition required by the generalized entropy accumulation theorem [MFSR22]. This allows us to lower bound the smooth min-entropy conditioned on the guesses and other side information of the modified adversary using the generalized entropy accumulation theorem. Finally, we derive an upper bound on $\Pr[\Omega]$ by using the fact that either the test fails with high probability, or the prover output has high entropy relative to the guesses and the guesser must fail. A rigorous proof is given in Supplement Section Appendix A.
@@ 4.3 Rapid-Firing
Construction 3.
In this construction, we assume the same syntax for $\mathscr{P}=(V,P)$ as in 2. In addition, we assume that $V$ samples challenges non-adaptively. That is, if $V=(\mathsf{Gen}_{1},\dots,\mathsf{Gen}_{\ell},\mathsf{Ver})$, then $\mathsf{Gen}_{i}$ ignores the inputs $\mathsf{ans}_{1},\dots,\mathsf{ans}_{i-1}$, that is, there exists $\widetilde{\mathsf{Gen}}_{i}$ such that
@@
Let $\mathscr{P}=(V,P)$ be a PoQ scheme that satisfies Definition 3, such that $V$ non-adaptively samples challenges. Then, 3 is a sound CVPV scheme in the quantum random oracle model.
The proof of Theorem 8 is nearly identical to the proof of Theorem 7. The major difference is that in the final hybrid we reach, the adversary has no time to communicate anymore due to the rapid-fire design, hence Definition 3 suffices. For the sake of completeness, we provide the full proof in Section C.1.
@@ 4.4 Sequential Rapid-Firing
Disclaimer
This paper was prepared for informational purposes by the Global Technology Applied Research center of JPMorgan Chase & Co. This paper is not a product of the Research Department of JPMorgan Chase & Co. or its affiliates. Neither JPMorgan Chase & Co. nor any of its affiliates makes any explicit or implied representation or warranty and none of them accept any liability in connection with this paper, including, without limitation, with respect to the completeness, accuracy, or reliability of the information contained herein and the potential legal, compliance, tax, or accounting effects thereof. This document is not intended as investment research or investment advice, or as a recommendation, offer, or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction.
Acknowledgments
We thank Dakshita Khurana and Kabir Tomer for pointing out an inaccuracy in our citation of recent work.
Appendix A CVPV from Certified Randomness through Repetition
@@
To proceed, we first provide a formal model of the sequential protocol quantum channels in Section A.1. We then consider an optimal adversary with the largest acceptance probability $\Pr[\Omega]$ in Section A.2. We also construct a modified adversary from this optimal adversary, and show that the modified adversary has the same $\Pr[\Omega]$. Then, we show the modified adversary also satisfies the non-signaling condition required by the generalized entropy accumulation theorem in Section A.3, which allows us to apply entropy accumulation to lower bound the entropy of the prover output conditioned on the guesser side information. Finally, we derive an upper bound on $\Pr[\Omega]$ by using the fact that either the certified randomness protocol test condition $\omega^{\prime}$ fails with high probability, or the prover output $A_{1}^{n}$ has high entropy relative to the guesser side information $E_{n}$ in Section A.4. This concludes the proof of the sequential decomposition property.
A.1 Protocol Quantum Channels
@@
$\mathcal{T}_{i}$ Quantum channel of the $i$th round prover test result computation. Explanation for the subscript $i$ will be omitted below
@@
$\Omega$ Channel to determine if the protocol passes from $C_{1}^{n}A_{1}^{n}G_{1}^{n}$, i.e. $\Omega=\Omega^{\prime}\circ\mathcal{T}_{n}\circ\dots\circ\mathcal{T}_{1}$. Also denotes the event of the success
@@
Table 1: Table of variables for this Appendix A.
@@
A.2 Proof of Equal Abort Probability
@@
where $\Gamma_{i}:C_{1}^{i}G_{1}^{i}\rightarrow R_{i}E_{i}$ is a quantum channel give by
@@ A.3 Proof of Entropy
A.4 Proof of the Sequential Decomposition Property
@@
Appendix B Instantiation

### v3

@@ 4 CVPV from Single-Round Certified Randomness
Completeness. Completeness follows by completeness of $\mathscr{P}$.
@@
Let $\mathscr{P}=(V,P)$ be a one-round PoQ scheme that satisfies Definition 3. Then, 6 is a CVPV scheme that is sound in the quantum random oracle model.
@@
  • The guesser $Q$ holds register ${\color[rgb]{1,.5,0}B}$. She receives $\mathsf{ch}$ and runs $\mathcal{B}(t\geq 0)$ to output $\mathsf{ans}^{\prime}$.
With probability $p_{4}$, $\mathsf{ans}=\mathsf{ans}^{\prime}$ and $\mathsf{ans}$ is accepted by the verifier, which violates certified randomness (Definition 3). ∎
@@ 5 CVPV from Multi-Round Certified Randomness
If the underlying PoQ-based certified randomness protocol uses more than one round, we can naturally generalize our compiler just by composing our single-round compiler sequentially. We first give a natural way to do this in Section 5.2. This requires a protocol satisfying the stronger Definition 6 of certified randomness with sequential decomposition, but we show that any protocol satisfying the definition of certified randomness with single round entropy must satisfy the sequential decomposition property [^2] In fact, we show that Definition 5 is satisfied which is stronger than Definition 6. in Section 7, lending its applicability to a wide range of possible instantiations. We then give a more clever way in Section 5.4 and Section 5.5 which is superior in the idealized model at the cost of practical robustness.
5.1 Multi-Round Certified Randomness: Definitions
We give three definitions of multi-round certified randomness, from strongest to weakest. While all definitions generalize Definition 3, they differ in terms of how much communication is allowed between the prover and the guesser in-between rounds: it is unrestricted (Definition 5), restricted (Definition 6), or forbidden (Definition 7).
@@ 5.1 Multi-Round Certified Randomness: Definitions
An $\ell$-round PoQ protocol $\mathscr{P}$ is said to have certified randomness with sequential decomposition property if no pair of a QPT prover $P$ and a (possibly entangled) unbounded guesser $Q$ can succeed in the following security game with non-negligible probability:
@@
We consider weaker variants of Definition 5 by restricting the communication between $P$ and $Q$ in-between rounds, with the differences being highlighted.
Below, we say that two spatially separated parties $(\mathcal{A},\mathcal{B})$ perform one round of simultaneous communication at time $t$ if $\mathcal{A}$ sends one (classical or quantum) message to $\mathcal{B}$ at time $t$ and vice versa.
@@
An $\ell$-round PoQ protocol $\mathscr{P}$ is said to have sequential certified randomness with sequential decomposition and restricted communication property if no pair of a QPT prover $P$ and a (possibly entangled) unbounded guesser $Q$ can succeed in the following security game with non-negligible probability:
@@
Definition 7 (Multi-Round Sequential Certified Randomness with Sequential Decomposition and No Communication).
An $\ell$-round PoQ protocol $\mathscr{P}$ is said to have sequential certified randomness with sequential decomposition and no communication property if no pair of a QPT prover $P$ and a (possibly entangled) unbounded guesser $Q$ can succeed in the following security game with non-negligible probability:
  • For $i\in[\ell]$, the following steps occur in order:
    1. 1.
    2. 2.
      $P$ sends back an answer $\mathsf{ans}_{i}$.
    3. 3.
      $Q$ outputs a guess $\mathsf{ans}_{i}^{\prime}$.
    4. 4.
      $P$ and $Q$ cannot communicate.
@@
For single-round $\mathscr{P}$, Definitions 3, 8, 5, 6 and 7 all coincide.
@@ 5.2 Sequential Compiler
Let $\mathscr{P}=(V,P)$ be a PoQ scheme that satisfies Definition 6. Then, 8 is a secure CVPV scheme.
@@
Follows directly from Definition 6. $\mathcal{A}$ plays the role of the prover and $\mathcal{B}$ that of the guesser. ∎
@@ 5.3 Sequential Decomposition from Repetition
This section uses the same notation as Section 7. Although other multi-round approaches are possible, the most general sequential compilation approach is desirable due to practical robustness considerations. Section 5.2 proves that one can construct CVPV protocols from certified randomness protocols with the sequential decomposition property. However, the usual notion of certified randomness (Definition 3) is defined in terms of entropy lower bounds, but we cannot lower bound the CVPV protocol entropy conditioned on the guesser’s side information since she can communicate with the prover. Therefore, it is not immediately obvious whether a certified randomness protocol defined with entropy is a sequential certified randomness protocol (Definition 5) defined by the guessing probability. We show that for a large class of multi-round certified randomness protocols, namely those that rely on repetition of single rounds with entropy bounds ‘on average’, all have the sequential decomposition property.
@@ A PoQ protocol $\mathscr{P}$ is said to be certified randomness from repetition if
To show that protocols satisfying Definition 8 can be used to construct CVPV protocols (illustrated in Fig. 7(b)) with the sequential decomposition property, we first consider an optimal adversary with the largest overall acceptance probability $\Pr[\Omega]$, where $\Omega$ denotes the event where the answers pass the protocol statistical test and all the guesses are correct. We denote the quantum channel of the $n$-round optimal adversary as $\mathcal{M}_{n}^{{}^{\prime}*}\circ\cdots\circ\mathcal{M}_{1}^{{}^{\prime}*}$. We then construct a modified adversary ¯$\bar{\mathcal{M}}_{n}^{{}^{\prime}}\circ\cdots\circ\bar{\mathcal{M}}_{1}^{{}^{\prime}}$ from this optimal adversary, and show that the modified adversary has the same $\Pr[\Omega]$. Then, we show the modified adversary also satisfies the non-signalling condition required by the generalized entropy accumulation theorem [MFSR22]. This allows us to lower bound the smooth min-entropy conditioned on the guesses and other side information of the modified adversary using the generalized entropy accumulation theorem. Finally, we derive an upper bound on $\Pr[\Omega]$ by using the fact that either the test fails with high probability, or the prover output has high entropy relative to the guesses and the guesser must fail. A rigorous proof is given in Section 7.
@@ 5.4 Rapid-Firing
Construction 10.
In this construction, we assume the same syntax for $\mathscr{P}=(V,P)$ as in 8. In addition, we assume that $V$ samples challenges non-adaptively. That is, if $V=(\mathsf{Gen}_{1},\dots,\mathsf{Gen}_{\ell},\mathsf{Ver})$, then $\mathsf{Gen}_{i}$ ignores the inputs $\mathsf{ans}_{1},\dots,\mathsf{ans}_{i-1}$, that is, there exists $\widetilde{\mathsf{Gen}}_{i}$ such that
@@
Let $\mathscr{P}=(V,P)$ be a PoQ scheme that satisfies Definition 7, such that $V$ non-adaptively samples challenges. Then, 10 is a sound CVPV scheme in the quantum random oracle model.
Proof.
The proof of Theorem 11 is nearly identical to the proof of Theorem 9. The major difference is that in the final hybrid we reach, the adversary has no time to communicate anymore due to the rapid-fire design, hence Definition 3 suffices. As in the proof of Theorem 9, we will consider hybrid experiments:
  • Hybrid 0: This is the original CVPV soundness experiment.
    with $u_{i}\leftarrow\{0,1\}^{n}$ being a fresh random string.
Claim 11.
$p_{1}\geq p_{0}$.
Proof.
Claim 12.
$p_{2}\geq p_{1}$.
Proof.
Claim 13.
Proof.
  • $\widetilde{\mathcal{B}}$ simulates $\mathcal{B}$.
Claim 14.
$p_{3}\geq p_{2.\ell}$.
Proof.
Claim 15.
$p_{3}\leq\mathsf{negl}(\lambda).$
Proof.
Follows from the certified randomness property (Definition 7). $\mathcal{A}$ plays the role of the prover and $\mathcal{B}$ that of the guesser. Note that the no-communication condition is satisfied because $(\ell-1)\Delta<1$, which means no message can be sent between $\mathcal{A}$ and $\mathcal{B}$ in time from the first challenge until the last. ∎
@@ 5.5 Sequential Rapid-Firing
6 Certified Randomness from CVPV
Recall that our work is partially motivated by discovering the fundamental property needed to construct CVPV. We have shown how to construct CVPV using certifed randomness, but so far have not discussed the reverse direction. In this section, we complete the picture by showing the reverse direction. To this end, we define a weaker version of certified randomness
6.1 Weak Certified Randomness: Definitions
We define a weak version of certified randomness by restricting the guesser to be computationally bounded.
Below, we give two definitions, one for single-round and one for multi-round. For the multi-round case, we use the weakest definition (Definition 10) which does not allow communication in-between rounds. Since this definition can be used to achieve CVPV, we do not expect We highlight deviations from the corresponding definitions (Definitions 3 and 7).
Definition 9 (Weak Certified Randomness (Single-Round)).
A proof of quantumness protocol $\mathscr{P}$ is said to have weak certified randomness property if no pair of a QPT prover $P$ and a QPT guesser $Q$ can succeed in the following security game with non-negligible probability:
  1. 1.
    The verifier $V$ of $\mathscr{P}$ sends a challenge $\mathsf{ch}$ to both $P$ and $Q$.
  2. 2.
    $P$ sends back an answer $\mathsf{ans}$ and $Q$ and outputs a guess $\mathsf{ans}^{\prime}$.
  3. 3.
    $(P,Q)$ win if $V$ accepts and $\mathsf{ans}=\mathsf{ans}^{\prime}$.
Definition 10 (Weak Multi-round Sequential Certified Randomness with Sequential Decomposition and No Communication).
An $\ell$-round PoQ protocol $\mathscr{P}$ is said to have weak sequential certified randomness with sequential decomposition and no communication property if no pair of a QPT prover $P$ and a QPT guesser $Q$ can succeed in the following security game with non-negligible probability:
  • For $i\in[\ell]$, the following steps occur in order:
    1. 1.
    2. 2.
      $P$ sends back an answer $\mathsf{ans}_{i}$.
    3. 3.
      $Q$ outputs a guess $\mathsf{ans}_{i}^{\prime}$.
    4. 4.
      $P$ and $Q$ cannot communicate.
Note that with this relaxation the definition still suffices for our CVPV construction since in the security proof, the guesser corresponds to a spoofer for position-verification, hence is computationally bounded. This holds true in both the single-round and the multi-round setting. We list the formal statements below for completeness:
Theorem 13 (Single-Round (V2)).
Let $\mathscr{P}=(V,P)$ be a one-round PoQ scheme that satisfies Definition 9. Then, 6 is a CVPV scheme that is sound in the quantum random oracle model.
Theorem 14 (Rapid-Firing (V2)).
Let $\mathscr{P}=(V,P)$ be a PoQ scheme that satisfies Definition 10, such that $V$ non-adaptively samples challenges. Then, 10 is a sound CVPV scheme in the quantum random oracle model.
6.2 Weak Certified Randomness from CVPV
Following the argument [LLQ22] that shows CVPV implies PoQ, we show that it implies weak certified randomness. Theorem 14 show that weak certified randomness with restricted communication (Definition 10) is the fundamental property required to construct CVPV [^4] To be precise, we also require non-adaptive challenge sampling for this no communication in-between rounds setting..
Theorem 15.
If there exists a CVPV scheme (Definition 4), then there exists a PoQ protocol with weak certified randomness (Definition 10).
Proof.
The construction is natural, and given in [LLQ22]. Without loss of generality, let $V_{0},V_{1}$ be the left, right verifiers and let $P^{*}$ be the honest prover of the CVPV scheme. We construct a PoQ protocol $\mathscr{P}=(V,P)$ as follows:
  • Verifier: $V$ runs $V_{0},V_{1}$ together, sending and expecting messages consistent with the timing constraints imposed by the CVPV scheme.
  • Prover: $P$ simply runs $P^{*}$, sending all messages to $V$.
In more detail, if $t=t_{1}$ is the time $P^{*}$ receives the first message(s) from $(V_{0},V_{1})$, then $V$ sends that message(s) as the first challenge $\mathsf{ch}_{1}$. Then, $P$ will send all messages $P^{*}$ is supposed to send before receiving any new information as $\mathsf{ans}_{1}$, and so on until the entire CVPV protocol is covered.
Onto the proof of security, suppose there exists an adversary $(P,Q)$ breaking the weak certified randomness with non-negligible probability $\epsilon$. We construct $(\mathcal{A},\mathcal{B})$ that breaks the CVPV security. $\mathcal{A}$ will be located at the position of $V_{0}$ and run $P$, whereas $\mathcal{B}$ will be located at the position of $V_{1}$ and run $Q$. In each round, after receiving the challenges from the nearest verifier, $\mathcal{A}$ and $\mathcal{B}$ will forward the challenges to each other. Once they receive both challenges, they will run $P,Q$ respectively and send the answer to the nearest verifier. With this reduction, it is easy to see that there is a perfect correspondence between the security game of the CVPV and weak certified randomness of $\mathscr{P}$. Therefore, $(\mathcal{A},\mathcal{B})$ beat the CVPV security with probability $\epsilon$, which is contradiction.
Remark 6.
Note that if we have an additional guarantee on the space-time structure of the CVPV scheme, it will yield a stronger notion of weak certified randomness. For instance, if the CVPV has a sequential structure so as to only allow simultaneous single-round communication in between challenges, then it would yield the weak version of Definition 6, which can be similarly defined. In this sense, our compilers are optimal in the random oracle model.
7 CVPV from Certified Randomness through Repetition
@@
To proceed, we first provide a formal model of the sequential protocol quantum channels in Section 7.1. We then consider an optimal adversary with the largest acceptance probability $\Pr[\Omega]$ in Section 7.2. We also construct a modified adversary from this optimal adversary, and show that the modified adversary has the same $\Pr[\Omega]$. Then, we show the modified adversary also satisfies the non-signaling condition required by the generalized entropy accumulation theorem in Section 7.3, which allows us to apply entropy accumulation to lower bound the entropy of the prover output conditioned on the guesser side information. Finally, we derive an upper bound on $\Pr[\Omega]$ by using the fact that either the certified randomness protocol test condition $\omega^{\prime}$ fails with high probability, or the prover output $A_{1}^{n}$ has high entropy relative to the guesser side information $E_{n}$ in Section 7.4. This concludes the proof of the sequential decomposition property.
7.1 Protocol Quantum Channels
@@
$\mathcal{T}_{i}$ Quantum channel of the $i$th round prover test result. Explanation for the subscript $i$ will be omitted below
@@
$\Omega$ $\Omega=\Omega^{\prime}\circ\mathcal{T}_{n}\circ\dots\circ\mathcal{T}_{1}$. Also denotes the event of the success
@@
Table 1: Table of variables for this Section 7.
@@
7.2 Proof of Equal Abort Probability
@@
where $\Gamma_{i}:C_{1}^{i}G_{1}^{i}\rightarrow R_{i}E_{i}$ is a quantum channel given by
@@ 7.3 Proof of Entropy
7.4 Proof of the Sequential Decomposition Property
@@
8 Instantiations
We give several instantiations for the certified randomness scheme used by our CVPV compiler. Our main instantiation (Section 8.1) is based on random circuit sampling, and utilizes our technical work in Section 7.
In addition, we give instantiations using error-correcting codes (Section 8.2) and noisy trapdoor claw-free functions (Section 8.3). These two sections do not require extra technical work, and they should mainly serve to demonstrate the generality of our compilers.
8.1 Instantiation Using Random circuit sampling